Chapter 3 Multitasking Maniacs Co. Multitasking Maniacs Co. manufactures and sells a specialized cell phone. The companys contribution format income statement for the most recent year is given below: %of Sales 100% Total Per Unit Sales (20,000 units) 1,200,000 60 Variable Expense Contribution Margin Fixed Expense Net Income 900,000 $45 $300,000 $15 $240,000 60,000 Management is anxious to increase the companys profit and has asked for an analysis of a number of items. Required 1. Compute the companys CM ratio 2. Compute the companys break-even point in both units and sales dollars 3. Compute the companys margin of safety in both units and sales dollars for the most recent af4. Assume that due to the addition of a new sales location, costing $62,000 per year in rent, sales increase by $400,000 next year. This change does not require any additional fixed cost as the employees required are simply being transferred from another sales location where business is slow. If cost behavior patterns remain unchanged (i.e., variable cost to produce and sell one cell phone stay the same), by how much will the companys net operating income increase 5. Refer to the original data. Assume that next year management wants the company to earn a profit of at least 5266,000 and sales projections show that they will sell 22,000 units. What would the sales price need to be in order to meet this target profit?
Answer #1

1. CM ratio = Contribution margin/Sales = $300000/$1200000 = 25%

2. Break-even point in units = Fixed expenses/Contribution per unit = $240000/$15 = 16000

Break-even point in sales dollars = 16000 units x $60 = $960000

OR

Fixed expenses/CM ratio = $240000/25% = $960000

3. Margin of safety (units) = Actual sales units - Break-even sales units = 20000 - 16000 = 4000

Margin of safety (sales dollars) = Actual sales - Break-even sales = $1200000 - $960000 = $240000

4. Increase in net operating income: $38000

Increase in contribution margin (25% x $400000) $       1,00,000
Less: Increase in fixed costs 62000
Increase in net operating income $           38,000

5. Sales price: $68 per unit

Profit = [(Sales price - Variable cost) x Quantity sold] - Fixed expense

Assuming the sales price to be 'X'

$266000 = [(X - $45) x 22000] - $240000

$266000 = 22000X - $990000 - $240000

$266000 + $990000 + $240000 = 22000X

$1496000 = 22000X

X = $1496000/22000 = $68